# MoEGCL: Mixture of Ego-Graphs Contrastive Representation Learning for Multi-View Clustering

### v1

@@ Introduction
With the rapid advancement of digitalization, data is increasingly being collected from multiple heterogeneous sources. For example, autonomous vehicles rely on multi-camera systems to capture diverse visual perspectives for decision-making. In structural biology, proteins exhibit complex quaternary structures that can be characterized through different analytical techniques. Modern medical diagnostics integrate multimodal data from various clinical tests to enhance accuracy. The term “multi-view data” is used to describe an item from multiple data sources. Multi-View Clustering (MVC)  (Tang:65) aims to integrate heterogeneous data from multiple views to discover meaningful cluster structures in an unsupervised manner, playing a pivotal role in modern data mining applications. These methods utilize a view-specific encoder network to generate a powerful representation, which is fused from various views for clustering tasks. However, there exists the issue of heterogeneity among data from different views. To solve the issue, many alignment methods have been proposed. For example, some methods use KL divergence to align the label or representation distributions among different views  (hershey2007approximating). Additionally, MVC employs contrastive learning (CL) to align representations across different views.
Figure 1: Mixture of Ego-Graphs Fusion (MoEGF). Firstly, construct an ego graph for each sample in each view. Secondly, input each ego graph as an expert into the Mixture-of-Experts (MoE) architecture and calculate the importance of each expert. Thirdly, merge the importance coefficient with the experts to form a fused ego graph. Lastly, aggregate the ego graphs of all samples to generate a global connected graph.
In recent years, Graph Neural Networks (GNNs) have emerged as a powerful paradigm for extracting and processing graph-structured information, achieving remarkable success across various domains. Although GNNs have significantly improved the MVC, the problem of coarse-grained graph fusion still exists. For instance, existing methods like DFMVC (ren:75) and CAGL (wang:76) first construct an independent graph for each view and then perform weighted fusion at the view level. However, this approach assigns a fixed fusion coefficient to the graph structure of each view, ignoring the differences between samples. This fusion approach operates at a coarse granularity and fails to capture fine-grained structural relationships, which ultimately limits performance in deep multi-view clustering tasks.
@@
There are mainly two kinds of methods in Multi-View Clustering (MVC) (Chem2022review): Shallow Multi-View Clustering methods and Deep Multi-View Clustering methods.
Shallow Multi-View Clustering methods (Chem2022review) are classified into two categories: Graph Multi-View Clustering and Subspace Multi-View Clustering. Graph Multi-View Clustering methods (fang2022structure; pan2021multi) generally follow these steps: construct view-specific graphs, followed by the creation of a fusion graph from several view-specific graphs using different regularization terms, and finally the generation of the clustering results using spectral clustering, graph-cut approaches, or other algorithms. Subspace Multi-View Clustering methods (Zhang:64; Huang:67) create a common subspace self-representation vector from every view, and then separate the samples into different subspaces using different regularization terms on the consensus self-representation matrix.
Many recent works have focused on Deep Multi-View Clustering, motivated by encouraging developments in deep learning (Yu:66; Xu:68). Specifically, these methods encode the non-linear feature using a deep neural network. The adversarial methodology is used in Deep Multi-View Clustering methods (li2021multi; wang2022adversarial) to coordinate patterns of hidden representations from different views and learn the latent representation. The attention mechanism is used by Zhou et al. (zhou2020end) to allocate each view a weight value. To obtain a consensus representation, they then use the weighted total of all view-wise presentations. The consensus representation is produced by Wang et al. (wang2022adversarial) using a weighted aggregation and $l_{1,2}$-norm restriction. Contrastive Learning (CL) may align representations from different perspectives at the sample level, which facilitates label distribution alignment. These CL methods (trosten2021reconsidering; xu2022multi) have outperformed the previous distribution alignment strategies in Multi-View Clustering, yielding state-of-the-art results on numerous public datasets.
@@
Jacobs et al. (jacobs:55) present the first Mixture-of-Experts (MoE) structure. It is viewed as a tree-structured framework that employs the divide-and-conquer tactic. Sparsely-gated MoE (Noam:77) is the first method to show significant gains in model capacity, training time, or model quality when gating is used. MH-MoE (wu:80) divides each input token into many smaller tokens, which are then assigned to and processed by a variety of experts concurrently before being smoothly reintegrated into the original token form.
We use a Mixture-of-Experts network to implement fine-grained graph fusion at the sample level. To the best of our knowledge, we are the first to present the fusion of ego graphs.
@@
We create the representation of each view by using the Autoencoder architecture (song2018self). It contains two parts: an encoder and a decoder. For the $m$-th view, $f^{m}$ represents the encoder module. The encoder module generates the representation of each view as follows:
@@
We use the Mixture-of-Experts Networks to integrate ego graphs. It consist of gateing network and an expert network. We concatenate the vectors of $M$ views as input to the gating network.
@@
We calculate the gating coefficients of experts related to input $z_{i}$. It is implemented through a multi-layer perceptron neural network $mlp^{1}$, followed by classification using a softmax activation function. The number of output categories by softmax is $M$. The specific calculation is as follows:
@@
$V_{i}$ denotes the fused adjacency vector of the $i$-th sample. Aggregating all adjacency vectors can generate an adjacency matrix.
@@
To improve the clustering results using both global features and structural data, we additionally use the Graph Convolutional Networks (GCN) (Kipf:74) to generate boosted graph representations. We adopt a two-layer GCN module to obtain the topological structure information. The graph representation $\tilde{Z}_{i}$ is computed as follows:
@@
Clustering Module
[…]
Experiments
Benchmark Datasets
Table 1: Summary of the six public datasets.
Datasets Samples Views Clusters #-view Dimensions
Caltech5V 1400 5 7 [40, 254, 1984, 512, 928]
WebKB 1051 2 2 [2949, 334]
LGG 267 4 3 [2000, 2000, 333, 209]
MNIST 60000 3 10 [342, 1024, 64]
RGBD 1449 2 13 [2048, 300]
LandUse 2100 3 21 [20, 59, 59]
Compared Methods
Implementation Details
@@
Evaluation Metrics
@@
Figure 4: The convergence analysis on the Caltech5V dataset. In the figure, the test ACC, NMI, and PUR are shown at the top, and the training loss is depicted at the bottom. It can be observed that around $400$ epochs, MoEGCL reaches a steady state, with the training loss no longer decreasing.
Figure 5: The parameter analysis on the Caltech5V dataset. The figure shows the changes in three evaluation metrics: ACC, NMI, and PUR. The metrics are influenced by two hyperparameters $\lambda$ and $\tau$. $\lambda$ is the combination coefficient of two loss functions. $\tau$ denotes temperature coefficient.
Experimental comparative results
@@
The convergence and generalization capabilities of MVC are crucial. To validate the convergence and generalization effectiveness of MoEGCL, we conduct tests on the Caltech5V dataset. The test ACC, NMI, PUR, and training loss are shown in Figure 4.
@@
This paper proposes a novel Mixture of Ego-Graphs Contrastive Representation Learning (MoEGCL) framework. It aims to implement fine-grained graph fusion for deep multi-view clustering tasks. It is composed of two modules, Mixture of Ego-Graphs Fusion (MoEGF) and Ego Graph Contrastive Learning (EGCL). The MoEGF module, which constructs ego graphs and utilizes the Mixture-of-Experts network, is presented for the fusion of ego graphs at the sample level. Furthermore, the EGCL module is proposed to lessen conflicts in the same clusters and get better semantic structure information. It further improves fine-grained graph representation. Extensive experiments demonstrate that MoEGCL achieves superior performance over state-of-the-art methods in deep multi-view clustering tasks. Specifically, we compare eight baseline methods on six public datasets to verify the effectiveness of our proposed method. In summary, MoEGCL is a simple and effective representation learning method. We believe it has good application prospects.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 22574146), the National Key Research and Development Program of China (Grant No. 2021ZD0201501), the Youth Foundation Project of Zhejiang Province (Grant No. LQ22F020035), the National Natural Science Foundation of China (No. 32200860), the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024SSYS0007), the Zhejiang provincial “Ten Thousand Talents Program” (2021R52007), the National Key R&D Program of China (2022YFB4500405).

### v2

@@ Introduction
With the rapid advancement of digitalization, data is increasingly being collected from multiple heterogeneous sources (zou2024look; zoudon; zhu:82). For example, autonomous vehicles rely on multi-camera systems to capture diverse visual perspectives for decision-making. In structural biology, proteins exhibit complex quaternary structures that can be characterized through different analytical techniques. Modern medical diagnostics integrate multimodal data from various clinical tests to enhance accuracy. The term “multi-view data” is used to describe an item from multiple data sources. Multi-View Clustering (MVC)  (Tang:65; zou2023inclusivity; xiao2024dual) aims to integrate heterogeneous data from multiple views to discover meaningful cluster structures in an unsupervised manner, playing a pivotal role in modern data mining applications. These methods utilize a view-specific encoder network to generate a powerful representation, which is fused from various views for clustering tasks. However, there exists the issue of heterogeneity among data from different views. Many alignment methods have been proposed to solve this issue. For example, some methods use KL divergence to align the label or representation distributions among different views  (hershey2007approximating; liusparsemvc; zou2023dpnet). Additionally, MVC employs contrastive learning (CL) to align representations across different views.
Figure 1: Mixture of Ego-Graphs Fusion. Firstly, an ego graph for each sample in each view is constructed, and input each ego graph as an expert into the Mixture-of-Experts (MoE) architecture and calculate the importance of each expert. Then, the importance coefficient with the experts are merged to form a fused ego graph. Lastly, aggregate the ego graphs of all samples to generate a global connected graph.
In recent years, Graph Neural Networks (GNNs) have emerged as a powerful paradigm for extracting and processing graph-structured information, achieving remarkable success across various domains. Although GNNs have significantly improved the MVC, the problem of coarse-grained graph fusion still exists. For instance, existing methods like DFMVC (ren:75) and CAGL (wang:76) first construct an independent graph for each view and then perform weighted fusion at the view level. However, this approach assigns a fixed fusion coefficient to the graph structure of each view, ignoring the differences between samples. This fusion approach operates at a coarse granularity and fails to capture fine-grained structural relationships, which ultimately limits the performance in deep MVC tasks.
@@
There are mainly two kinds of methods in Multi-View Clustering (MVC) (Chem2022review; zhu2025generative): Shallow Multi-View Clustering methods and Deep Multi-View Clustering methods. Shallow MVC methods (Chem2022review) are classified into two categories: Graph Multi-View Clustering and Subspace Multi-View Clustering. Graph Multi-View Clustering methods (fang2022structure; pan2021multi) generally follow these steps: construct view-specific graphs, followed by the creation of a fusion graph from several view-specific graphs using different regularization terms, and finally the generation of the clustering results using spectral clustering, graph-cut approaches, or other algorithms. Subspace Multi-View Clustering methods (Zhang:64; Huang:67) create a common subspace self-representation vector from every view, and then separate the samples into different subspaces using different regularization terms on the consensus self-representation matrix.
Many recent works have focused on Deep MVC, motivated by encouraging developments in deep learning (zhu:81; zhu:83; zhu2025trusted; xu2025hstrans). Specifically, these methods encode the non-linear feature using a deep neural network. The adversarial methodology is used in Deep MVC methods (li2021multi; wang2022adversarial) to coordinate patterns of hidden representations from different views and learn the latent representation. The attention mechanism is used by zhou2020end; zhu2025dynamic to allocate each view a weight value. To obtain a consensus representation, they then use the weighted total of all view-wise presentations. The consensus representation is produced by Wang et al. (wang2022adversarial) using a weighted aggregation and $l_{1,2}$-norm restriction. Contrastive Learning (CL) may align representations from different perspectives at the sample level, which facilitates label distribution alignment. These CL methods (trosten2021reconsidering; xu2022multi) have outperformed the previous distribution alignment strategies in Multi-View Clustering, yielding state-of-the-art results on numerous public datasets.
@@
Jacobs et al. (jacobs:55) present the first Mixture-of-Experts (MoE) structure. It is viewed as a tree-structured framework that employs the divide-and-conquer tactic. Sparsely-gated MoE (Noam:77) is the first method to show significant gains in model capacity, training time, or model quality when gating is used. MH-MoE (wu:80) divides each input token into many smaller tokens, which are then assigned to and processed by a variety of experts concurrently before being smoothly reintegrated into the original token form. We use a Mixture-of-Experts network to implement fine-grained graph fusion at the sample level. To the best of our knowledge, we are the first to present the multi-view fusion framework of ego graphs.
@@
We create the representation of each view by using the Autoencoder architecture (song2018self; zou2023hierarchical; zou2024dai). It contains two parts: an encoder and a decoder. For the $m$-th view, $f^{m}$ represents the encoder. The encoder module generates the representation of each view as follows:
@@
We use the Mixture-of-Experts Networks to integrate ego graphs. It consists of a gating network and an expert network. We concatenate the vectors of $M$ views as input to the gating network.
@@
We calculate the gating coefficient for each expert with respected to input $z_{i}$. It is implemented through a multi-layer perceptron neural network $mlp^{1}$, followed by classification using a softmax activation function. The number of output categories by softmax is $M$. The specific calculation is as follows:
@@
$V_{i}$ denotes the fused adjacency vector of the $i$-th sample. Aggregating all adjacency vectors can generate an adjacency matrix,
@@
To improve the clustering results using both global features and structural data, we additionally use the Graph Convolutional Networks (GCN) (Kipf:74; zhang2019graph) to generate boosted graph representations. We adopt a two-layer GCN module to obtain the topological structure information. The graph representation $\tilde{Z}_{i}$ is computed as follows:
@@
Datasets Samples Views Clusters #-view Dimensions
Caltech5V 1400 5 7 [40, 254, 1984, 512, 928]
WebKB 1051 2 2 [2949, 334]
LGG 267 4 3 [2000, 2000, 333, 209]
MNIST 60000 3 10 [342, 1024, 64]
RGBD 1449 2 13 [2048, 300]
LandUse 2100 3 21 [20, 59, 59]
Table 1: Summary of the six public datasets.
@@
Clustering Module
[…]
Experiments
Benchmark Datasets
@@
Compared Methods
Implementation Details
Evaluation Metrics
Experimental comparative results
@@
Figure 4: The convergence analysis on the Caltech5V dataset. In the figure, the test ACC, NMI, and PUR are shown at the top, and the training loss is depicted at the bottom. It can be observed that around $400$ epochs, MoEGCL reaches a steady state, with the training loss no longer decreasing, and achieves the best accuracy compared with SOTA methods.
Figure 5: The parameter analysis on the Caltech5V dataset. The figure shows the changes in three evaluation metrics: ACC, NMI, and PUR. The metrics are influenced by two hyperparameters $\lambda$ and $\tau$. $\lambda$ is the combination coefficient of two loss functions. $\tau$ denotes the temperature coefficient.
@@
The convergence and generalization capabilities of MVC are crucial. To validate the convergence and generalization effectiveness of MoEGCL, we conduct tests on the Caltech5V dataset. The performance and loss are shown in Figure 4.
@@
This paper proposes a novel Mixture of Ego-Graphs Contrastive Representation Learning (MoEGCL) framework. It aims to implement fine-grained graph fusion for deep multi-view clustering tasks. It is composed of two modules, Mixture of Ego-Graphs Fusion (MoEGF) and Ego Graph Contrastive Learning (EGCL). The MoEGF module, which constructs ego graphs and utilizes the Mixture-of-Experts network, is presented for the fusion of ego graphs at the sample level. Furthermore, the EGCL module is proposed to lessen conflicts in the same clusters and get better semantic structure information. It further improves fine-grained graph representation. Extensive experiments demonstrate that MoEGCL achieves superior performance over state-of-the-art methods in deep multi-view clustering tasks. Specifically, we compare eight baseline methods on six public datasets to verify the effectiveness of our proposed method. In summary, MoEGCL is a simple and effective representation learning method, it has great application prospects.
@@ Acknowledgments
This work is supported by the National Key Research and Development Program of China (Grant No. 2021ZD0201501), the National Natural Science Foundation of China (No. 22574146, No. 32200860, and No. 62306289), by the Regional Innovation and Development Joint Fund of National Natural Science Foundation of China (U22A6001), the National Key R&D Program of China (2022YFB4500405), the Youth Foundation Project of Zhejiang Province (Grant No. LQ22F020035), the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024SSYS0007), the National Key R&D Program of China (Grant No. 2024YFB4505602), by Anhui Provincial Natural Science Foundation (Grant No.2408085QF214), the Fundamental Research Funds for the Central Universities (Grant No.WK2100000045, No.WK2080000206), and the Opening Project of the State Key Laboratory of General Artificial Intelligence (Grant No.SKLAGI2024OP10, Grant No.SKLAGI2024OP11).